OpenStudy (uuioo):

Two equations are given below: a − 3b = 4 a = b − 2 What is the solution to the set of equations in the form (a, b)?

OpenStudy (mathmale):

There are several different methods applicable here. You could use substitution, elimination by addition / subtraction, graphing. Which of these methods is most familiar to you? Try applying it to the problem at hand.

OpenStudy (mhchen):

since a=b-2 plug that into the first equation: (b-2) - 3b = 4 b-3b = 6 -2b = 6 b = -3 Then plug b into any equation: a = b - 2 a = (-3) - 2 a = -5

OpenStudy (mathmale):

mhchen is proposing that y ou use the method of substitution to solve this system of linear equations. Can you do that?